# Quantum Mixing and Benjamini–Schramm Convergence of Hyperbolic Surfaces

### v1

@@ 7. Weight function
The proof is based on using the properties of the isometries in the hyperbolic plane.
@@
There exists an isometry $f$ such that $f(z)$ equals $w$ and that $f(z^{\prime})$ equals $w^{\prime}$. Using a change of variables $x\mapsto f^{-1}(x)$, we get the following equal expression for the left-hand side of Expression (7.1):
@@
where the equality followed from the fact that $d(f^{-1}(x),z)=d(x,f(z))$ which then equals $d(x,w)$, and similarly $d(f^{-1}(x),z^{\prime})=d(x,w^{\prime})$. As $f$ maps $z$ to $w$ and $z^{\prime}$ to $w^{\prime}$, we have that
@@
The main goal of this section is to give an upper bound for $F_{t,t^{\prime},\rho}$. Here, we additionally assume that $t>t^{\prime}$ for the case $t^{\prime}<t$ is symmetric. $F_{t,t^{\prime},\rho}$ is a crucial weight appearing in Section 6 when proving bounds for the geometric data appearing in the proof of Theorem 1.5 in Proposition 6.1. Bounding this weight is done in three lemmas, Lemmas 7.3, 7.4, and 7.5, of which settings are illustrated, respectively, in the following three figures:
Figure 2. The case of $t>t^{\prime}+\rho$. The integration domain $B(z,t)\cap B(z^{\prime},t^{\prime})$ is just the smaller ball $B(z^{\prime},t^{\prime})$ for it is enveloped by the bigger ball $B(z,t)$. The bound in this case is established using the polar coordinates around $z^{\prime}$.
Figure 3. The case of $\rho<t<t^{\prime}+\rho$, where $a+b=\rho$. The integration domain $B(z,t)\cap B(z^{\prime},t^{\prime})$ is the intersection of two balls. The middle point $z^{\prime}$ of the smaller ball is contained in the bigger ball. The integration domain is split into two parts, $D_{1}$ and $D_{2}$ by two geodesics of length $t^{\prime}$ going from $z^{\prime}$ to points $w$ and $w^{\prime}$ which are the points where the two balls intersect. The bound in this case is established separately for $D_{1}$ and $D_{2}$. For $D_{1}$, we use the polar coordinates around $z^{\prime}$ and for $D_{2}$ the polar coordinates around $z$.
Figure 4. The case of $t<\rho$. The integration domain $B(z,t)\cap B(z^{\prime},t^{\prime})$ is the intersection of two balls. The middle point $z^{\prime}$ of the smaller ball is not contained in the bigger ball. The bound in this case is established using the polar coordinates around $m$ which is the intersection of geodesics $z,z^{\prime}]$ and $w,w^{\prime}]$, where $w$ and $w^{\prime}$ are the intersections of the two balls.
From these three lemmas we get the following bound for $F_{t,t^{\prime},\rho}$.
@@
$F_{t,t^{\prime},\rho}\leq C_{1}\,\cdot\,\begin{cases}\frac{e^{\frac{t^{\prime}-t}{4}}}{(\sinh(\rho)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}},\qquad&\text{ if }\rho<t-t^{\prime}\\ \frac{1}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{\frac{1}{8}},\qquad&\text{ if }t-t^{\prime}<\rho<t\\ \frac{1}{\sqrt{\sinh(\rho)}},\qquad&\text{ if }t<\rho<t^{\prime}+t.\end{cases}$
where
$C_{1}=\max\left(C_{2},C_{3},C_{4}\right)=\max(32,32\pi,99)=32\pi.$
The following lemma gives the bound for $F_{t,t^{\prime},\rho}$ in the first case, see Figure 2.
@@
Assume that $t$ is larger than $t^{\prime}+\rho$. Then
$F_{t,t^{\prime},\rho}\leq\frac{C_{2}\,e^{\frac{t^{\prime}-t}{4}}}{(\sinh(\rho)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}},$
where
$C_{2}=32.$
The proof of lemma is based on using two different lower bounds for $\cosh(t)-\cosh(R)$, when working in the polar coordinates around $z^{\prime}$. Using these two bounds simultaneously allows us to deal with the problematic small angles, when the boundaries of two balls are close to each other.
@@
We write the integral using the polar coordinates around $z^{\prime}$, taking $\theta=0$ in the direction $z,z^{\prime}](z^{\prime})$. Hence $x=\pi_{1}(\phi_{r}(z^{\prime},\theta))$. The situation is illustrated in Figure 2. Hence $F_{t,t^{\prime},\rho}$ is written as
$\int_{0}^{t^{\prime}}\int_{-\pi}^{\pi}\frac{\sinh(r)\,d\theta\,dr}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}$
$=\qquad$ $2\int_{0}^{t^{\prime}}\int_{0}^{\pi}\frac{\sinh(r)\,d\theta\,dr}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}},$ (7.2)
where half of the angles are dropped using symmetry, and where $R=R(r,\theta)$ is $d(x,z)$ using the polar coordinates $r$ and $\theta$.
We want to find a lower bound for $\cosh(t)-\cosh(R)$. We remark that we have the following inequality since $t>t^{\prime}+\rho>R$:
$\cosh(t)-\cosh(R)\geq\cosh(t)-\cosh(t^{\prime}+\rho).$
This implies the following inequality by using the mean value theorem and the convexity of the hyperbolic sine for the positive input:
$\cosh(t)-\cosh(R)\geq\sinh(t^{\prime})\,\sinh(\rho)\,(t-t^{\prime}-\rho).$ (7.3)
We can also get another inequality. We write $\cosh(t)$ as follows:
$\cosh(t)=\cosh(t-t^{\prime})\cosh(t^{\prime})+\sinh(t-t^{\prime})\sinh(t^{\prime}).$
Then, by using the hyperbolic law of cosines, we have that
$\cosh(R)=\cosh(\rho)\cosh(r)+\sinh(\rho)\sinh(r)\cos(\theta).$
Hence $\cosh(t)-\cosh(R)$ equals
$\cosh(t-t^{\prime})\cosh(t^{\prime})+\sinh(t-t^{\prime})\sinh(t^{\prime})-\cosh(\rho)\cosh(r)-\sinh(\rho)\sinh(r)\cos(\theta).$
Since $r<t^{\prime}$ and $\rho<t-t^{\prime}$, we have the following inequality:
$\cosh(t)-\cosh(R(r,\theta))\geq(1-\cos(\theta))\,\sinh(t)\,\sinh(t-t^{\prime}).$ (7.4)
We write Expression (7.2) as
$2\int_{0}^{t^{\prime}}\frac{\sinh(r)}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\int_{0}^{\pi}\frac{1}{(\cosh(t)-\cosh(R))^{\frac{1}{4}}\,(\cosh(t)-\cosh(R))^{\frac{1}{4}}}d\theta dr,$
which has the following upper bound
$\frac{2}{(\sinh(\rho)\,\sinh(t^{\prime})\,\sinh(t)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}}\int_{0}^{t^{\prime}}\frac{\sinh(r)}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\int_{0}^{\pi}\frac{1}{(1-\cos(\theta))^{\frac{1}{4}}}d\theta dr,$
which was acquired by using the inequalities appearing in Expressions (7.3) and (7.4). The following equal expressions are acquired by evaluating the integrals:
@@
$\frac{2}{(\sinh(\rho)\,\sinh(t^{\prime})\,\sinh(t)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}}\int_{0}^{t^{\prime}}\frac{\sinh(r)}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\,2^{\frac{5}{4}}\,K\left(\frac{1}{2}\right)dr$
@@
$=\ \frac{2^{2+\frac{5}{4}}\,K(\frac{1}{2})\,\sqrt{\cosh(t^{\prime}-1)}}{(\sinh(\rho)\,\sinh(t^{\prime})\,\sinh(t)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}},$
where $K$ is the complete elliptic integral of the first kind (see [Olv10, Chp.19]). The previous expression has a further upper bound, where we also bound the elliptic integral by $2$:
$\frac{2^{\frac{17}{4}}\,e^{\frac{t^{\prime}-t}{4}}}{(\sinh(\rho)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}}\leq 32\,\frac{e^{\frac{t^{\prime}-t}{4}}}{(\sinh(\rho)\,\sinh(t-t^{\prime})\,(t-t^{\prime}-\rho))^{\frac{1}{4}}}.$
We proceed to a lemma bounding the second case, see Figure 3.
@@
Assume that $\rho<t<\rho+t^{\prime}$. Then:
$F_{t,t^{\prime},\rho}\leq C_{3}\,\frac{1}{\sqrt{\sinh(\rho)}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}},$
where
$C_{3}=32\pi.$
The idea of the proof is to split the integration domain into two separate domains. The bounding of the other domain resembles the approach of the previous lemma in that we establish two different lower bounds for $\cosh(t)-\cosh(R)$, which allow us to get a bound when used simultaneously. Bounding the other domain turns out to be simple. The following figure, Figure 5, further illustrates these separate domains:
((a)) The integration domain $D_{1}$ consists of those points that are outside the triangle formed in the ball $B(z^{\prime},t^{\prime})$. In this case, the polar coordinates around $z^{\prime}$ are used.
((b)) The integration domain $D_{2}$ consists of those points that are inside the triangle formed in the ball $B(z^{\prime},t^{\prime})$. In this case, polar coordinates around $z$ are used.
Figure 5. The case is split into two case which are bounded separately.
@@
The integration domain of $F_{t,t^{\prime},\rho}$ can be split into a sum of integration over two disjoint domains, $D_{1}$ and $D_{2}$, see Figure 3. We bound these separately. $D_{1}$ is better illustrated in Figure 5(a) and $D_{2}$ in Figure 5(b).
Focus on $D_{1}$ first. We write it using the polar coordinates around $z^{\prime}$. Thus $x=\pi_{1}(\phi_{r}(z^{\prime},\theta))$, where $\theta=0$ corresponds to direction $z^{\prime},z](z^{\prime})$. In this way, the integral over $D_{1}$ equals
$\int_{-\Theta}^{\Theta}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}$
= $=\ 2\int_{0}^{\Theta}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}$
where half the angles were dropped using symmetry, and where $R=R(r,\theta)$ is $d(x,z)$ in the polar coordinates $r$ and $\theta$. Notice that the maximum angle $\Theta$ is larger than $\pi/2$. We split the integral into two terms as follows:
$2\int_{0}^{\frac{\pi}{2}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}$
+ $+\ 2\int_{\frac{\pi}{2}}^{\Theta}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr\,d\theta}{\sqrt{\cosh(t)-\cosh(R)}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}.$ (7.5)
By using the hyperbolic law of cosines we get that
$\cosh(t)=\cosh(\rho)\cosh(t^{\prime})-\sinh(\rho)\sinh(t^{\prime})\cos(\Theta),$ (7.6)
$\cosh(R)=\cosh(\rho)\cosh(r)-\sinh(\rho)\sinh(r)\cos(\theta).$ (7.7)
Hence $\cosh(t)-\cosh(R)$ equals
$\cosh(\rho)(\cosh(t^{\prime})-\cosh(r))+\sinh(\rho)(\cos(\theta)\sinh(r)-\cos(\Theta)\sinh(t^{\prime})).$ (7.8)
Hence we have the following inequalities:
$\cosh(t)-\cosh(R)\geq\sinh(\rho)\sinh(t^{\prime})(\cos(\theta)-\cos(\Theta))$ (7.9)
and
$\cosh(t)-\cosh(R)\geq\cosh(\rho)\,(\cosh(t^{\prime})-\cosh(r)).$ (7.10)
Using the inequality of Expression (7.9), we get that the first term of Expression (7.5) is bounded by
$2\int_{0}^{\frac{\pi}{2}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr\,d\theta}{\sqrt{\sinh(\rho)\sinh(t^{\prime})(\cos(\theta)-\cos(\Theta))}\,\sqrt{\cosh(t^{\prime})-\cosh(r)}}.$
By evaluating the integral over $r$, we get the following equal expression:
$\frac{4}{\sqrt{\sinh(\rho)}}\,\frac{\sqrt{\cosh(t^{\prime})-1}}{\sqrt{\sinh(t^{\prime})}}\int_{0}^{\frac{\pi}{2}}\frac{1}{\sqrt{\cos(\theta)-\cos(\Theta)}}\,d\theta.$
This has the following upper bound, when we notice that both $\cos(\theta)$ and $\cos(\Theta)$ are positive. Additionally the ratio with the hyperbolic sine and cosine is bounded by $1$:
$\frac{4}{\sqrt{\sinh(\rho)}}\,\int_{0}^{\frac{\pi}{2}}\frac{1}{\sqrt{\cos(\theta)}}\,d\theta.$
The following bound follows from the well known upper bound for the integral:
$\frac{12}{\sqrt{\sinh(\rho)}}.$
Let us focus on the second term of Expression (7.5). We write it as follows by interchanging the integrals using Fubini’s theorem:
$2\,\int_{0}^{t^{\prime}}\int_{\frac{\pi}{2}}^{\Theta}\frac{\sinh(r)}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\,\frac{1}{(\cosh(t)-\cosh(R))^{\frac{1}{4}}\,(\cosh(t)-\cosh(R))^{\frac{1}{4}}}\,d\theta\,dr.$
We use the two inequalities appearing in Expressions (7.9) and (7.10) to get the following upper bound for the previous expression:
$\frac{2}{\cosh(\rho)^{\frac{1}{4}}\,\sinh(\rho)^{\frac{1}{4}}\,\sinh(t^{\prime})^{\frac{1}{4}}}\,\int_{0}^{t^{\prime}}\frac{\sinh(r)}{(\cosh(t^{\prime})-\cosh(r))^{\frac{3}{4}}}\int_{\frac{\pi}{2}}^{\Theta}\frac{1}{(\cos(\theta)-\cos(\Theta))^{\frac{1}{4}}}\,d\theta\,dr.$
Using the mean value theorem, we get the following upper bound:
$\frac{2}{\cosh(\rho)^{\frac{1}{4}}\,\sinh(\rho)^{\frac{1}{4}}\,\sinh(t^{\prime})^{\frac{1}{4}}}\,\int_{0}^{t^{\prime}}\frac{\sinh(r)}{(\cosh(t^{\prime})-\cosh(r))^{\frac{3}{4}}}\,\frac{1}{\sin(\Theta)^{\frac{1}{4}}}\int_{\frac{\pi}{2}}^{\Theta}\frac{1}{(\Theta-\theta)^{\frac{1}{4}}}d\theta dr.$
Integrating the inner integral yields
$\frac{4}{3}\,\left(\Theta-\frac{\pi}{2}\right)^{\frac{3}{4}}.$
Hence we have the following bound for the previous integral expression:
$\frac{4\pi}{\cosh(\rho)^{\frac{1}{4}}\,\sinh(\rho)^{\frac{1}{4}}\,\sinh(t^{\prime})^{\frac{1}{4}}}\,\int_{0}^{t^{\prime}}\frac{\sinh(r)}{(\cosh(t^{\prime})-\cosh(r))^{\frac{3}{4}}}\,\frac{1}{\sin(\Theta)^{\frac{1}{4}}}dr$
$=\ \frac{16\pi}{\cosh(\rho)^{\frac{1}{4}}\,\sinh(\rho)^{\frac{1}{4}}\,\sinh(t^{\prime})^{\frac{1}{4}}}\,(\cosh(t^{\prime})-1)^{\frac{1}{4}}\,\frac{1}{\sin(\Theta)^{\frac{1}{4}}},$ (7.11)
where the equality followed from the evaluation of the remaining integral.
Focus momentarily on $\sin(\Theta)$. Expression (7.6) can be modified into the following form using the hyperbolic identities:
$\cos(\Theta)=1-\frac{\cosh(t)-\cosh(\rho-t^{\prime})}{\sinh(t^{\prime})\sinh(\rho)}.$
The subtracted division term must be larger than $1$ as the cosine is negative due to $\Theta$ being in $(\frac{\pi}{2},\pi)$. We have that
$\sin(\arccos(1-x))=\sqrt{x}\sqrt{2-x}\geq\sqrt{2-x},$
when $x>1$. Hence:
$\sin(\Theta)\geq\sqrt{2-\frac{\cosh(t)-\cosh(\rho-t^{\prime})}{\sinh(t^{\prime})\sinh(\rho)}}=\sqrt{\frac{\cosh(t^{\prime}+\rho)-\cosh(t)}{\sinh(t^{\prime})\sinh(\rho)}},$
where the equality followed using hyperbolic identities. The previous implies the following inequality:
$\sin(\Theta)\geq\sqrt{\frac{\cosh(t^{\prime}+\rho)-\cosh(t)}{\cosh(t^{\prime}+\rho)}}=\sqrt{1-\frac{\cosh(t)}{\cosh(t^{\prime}+\rho)}}.$
This then gives us the following inequality:
$\sin(\Theta)\geq\sqrt{1-\frac{1}{\cosh(t^{\prime}+\rho-t)}}.$ (7.12)
Applying the inequality of Expression (7.12) in Expression (7.11) gives the following upper bound:
$\frac{16\pi}{\cosh(\rho)^{\frac{1}{4}}\,\sinh(\rho)^{\frac{1}{4}}\,\sinh(t^{\prime})^{\frac{1}{4}}}\,(\cosh(t^{\prime})-1)^{\frac{1}{4}}\,\left(1-\frac{1}{\cosh(t^{\prime}+\rho-t)}\right)^{-\frac{1}{8}}.$
This has the following further upper bound:

### v2

@@ 7. Weight function
The proof is based on the properties of isometries in the hyperbolic plane.
@@
There exists an isometry $f$ such that $f(z)=w$ and $f(z^{\prime})=w^{\prime}$. Using a change of variables $x\mapsto f^{-1}(x)$, we get the following equal expression for the left-hand side of Expression (7.1):
@@
where the equality follows from the fact that $d(f^{-1}(x),z)=d(x,f(z))$, which then equals $d(x,w)$, and similarly $d(f^{-1}(x),z^{\prime})=d(x,w^{\prime})$. As $f$ maps $z$ to $w$ and $z^{\prime}$ to $w^{\prime}$, we have that
@@
The main goal of this section is to give an upper bound for $F_{t,t^{\prime},\rho}$. $F_{t,t^{\prime},\rho}$ is a crucial weight appearing in Section 6 when proving bounds for the geometric data of the main problem. The upper bound is given in the following proposition.
@@
Let
where $z,z^{\prime}\in\mathbb{H}$ are any two points separated by distance $\rho$. Assume that $t>t^{\prime}>0$.
$F_{t,t^{\prime},\rho}\leq C_{1}\,\frac{1}{\sqrt{\sinh(\max(t-t^{\prime},\rho))}}\bm{1}_{\rho\in(0,t+t^{\prime})},$
where $C_{1}=50$.
Remark 7.3.
The case of $t>t^{\prime}>0$ is symmetric to the case of $t^{\prime}>t>0$.
Despite the simplicity of the upper bound, obtaining it is tedious. The proof is based on carefully using hyperbolic geometry. In case $\rho\in(-\infty,0)\cup(t+t^{\prime},\infty)$, $F_{t,t^{\prime},\rho}=0$ since the integration domain is an empty set. The remaining data for the bound is obtained in three separate lemmas, Lemmas 7.4, 7.5, and 7.6, as the settings differ significantly from case to case. The remainder of the section is dedicated to stating and proving these three lemmas.
We state the first lemma concerning the case of $\rho\in(0,t-t^{\prime})$.
@@
Let
where $z,z^{\prime}\in\mathbb{H}$ are any two points separated by distance $\rho$. Assume that $t>t^{\prime}>0$ and that $\rho\in(0,t-t^{\prime})$. Then
$F_{t,t^{\prime},\rho}\leq\frac{C_{2}}{\sqrt{\sinh(t-t^{\prime})}},$
where $C_{2}=40$.
The following figure, Figure 2, illustrates the integration domain.
Figure 2. The case of $\rho\in(0,t-t^{\prime})$. The integration domain $B(z,t)\cap B(z^{\prime},t^{\prime})$ is just the smaller ball $B(z^{\prime},t^{\prime})$ since it is contained within the larger ball $B(z,t)$.
In the proof, polar coordinates around $z^{\prime}$ are used to present the integration domain $B(z,t)\cap B(z^{\prime},t^{\prime})=B(z^{\prime},t^{\prime})$. The main idea is to carefully split the integration into an integration over small angles and an integration over large angles; the small angles are especially delicate, as here $\partial B(z,t)$ and $\partial B(z^{\prime},t^{\prime})$ can be problematically close.
@@
We have that $B(z,t)\cap B(z^{\prime},t^{\prime})=B(z^{\prime},t^{\prime})$. Hence we write $x$ using polar coordinates around $z^{\prime}$, taking $\theta=0$ in the direction $z,z^{\prime}](z^{\prime})$. Hence $x=\pi_{1}(\varphi_{\rho}(z^{\prime},\theta))$. $R(r,\theta)$ is $d(x,z)$ in the polar coordinates. Thus:
@@
$F_{t,t^{\prime},\rho}=2\,\int_{0}^{t^{\prime}}\frac{\sinh(r)}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\left(\int_{0}^{\pi}\frac{d\theta}{\sqrt{\cosh(t)-\cosh(R(r,\theta))}}\right)dr,$ (7.2)
where only half of the angles are considered due to symmetry, thus giving the coefficient $2$ in front. We notice that by the hyperbolic law of cosines
$\cosh(R(r,\theta))=\cosh(r)\cosh(\rho)+\sinh(r)\sinh(\rho)\cos(\theta).$ (7.3)
We also remark that
$\cosh(t)=\cosh(t^{\prime})\cosh(t-t^{\prime})+\sinh(t^{\prime})\sinh(t-t^{\prime}).$ (7.4)
Focus on the inner integral of Expression (7.2). First, write it using Expressions (7.3) and (7.4):
$\int_{0}^{\pi}\frac{d\theta}{\sqrt{\cosh(t-t^{\prime})\cosh(t^{\prime})+\sinh(t-t^{\prime})\sinh(t^{\prime})-\cosh(r)\cosh(\rho)-\sinh(r)\sinh(\rho)\cos(\theta)}}.$
We split it into three terms by splitting the integration domain:
$\int_{0}^{\theta_{0}}\frac{d\theta}{\sqrt{\cosh(t-t^{\prime})\cosh(t^{\prime})+\sinh(t-t^{\prime})\sinh(t^{\prime})-\cosh(r)\cosh(\rho)-\sinh(r)\sinh(\rho)\cos(\theta)}}$
$+\ \int_{\theta_{0}}^{\frac{\pi}{2}}\frac{d\theta}{\sqrt{\cosh(t-t^{\prime})\cosh(t^{\prime})+\sinh(t-t^{\prime})\sinh(t^{\prime})-\cosh(r)\cosh(\rho)-\sinh(r)\sinh(\rho)\cos(\theta)}}$
$+\ \int_{\frac{\pi}{2}}^{\pi}\frac{d\theta}{\sqrt{\cosh(t-t^{\prime})\cosh(t^{\prime})+\sinh(t-t^{\prime})\sinh(t^{\prime})-\cosh(r)\cosh(\rho)-\sinh(r)\sinh(\rho)\cos(\theta)}}.$ (7.5)
$\theta_{0}\in(0,\frac{\pi}{2})$ is specified later; in particular, $\theta_{0}$ may depend on $r$.
We bound the first term of Expression (7.5). Since $t-t^{\prime}>\rho$ and $t^{\prime}>r$, we have the following upper bound for it:
$\int_{0}^{\theta_{0}}\frac{d\theta}{\sqrt{\cosh(t-t^{\prime})\cosh(t^{\prime})-\cosh(r)\cosh(\rho)}}=\frac{\theta_{0}}{\sqrt{\cosh(t-t^{\prime})\cosh(t^{\prime})-\cosh(r)\cosh(\rho)}}$
$\leq\ \frac{\theta_{0}}{\sqrt{\cosh(t-t^{\prime})}\sqrt{\cosh(t^{\prime})-\cosh(r)}},$
where the inequality follows from $t-t^{\prime}>\rho$.
Let us bound the second term of Expression (7.5). Since $t-t^{\prime}>\rho$ and $t^{\prime}>r$, we have the following upper bound for it:
$\int_{\theta_{0}}^{\frac{\pi}{2}}\frac{d\theta}{\sqrt{\sinh(t-t^{\prime})\sinh(t^{\prime})-\sinh(r)\sinh(\rho)\cos(\theta)}}\leq\frac{1}{\sqrt{\sinh(t-t^{\prime})\,\sinh(t^{\prime})}}\int_{\theta_{0}}^{\frac{\pi}{2}}\frac{d\theta}{\sqrt{1-\cos(\theta)}},$
where the inequality follows from $t-t^{\prime}>\rho$ and $t^{\prime}>r$. We have that
$1-\cos(x)\geq\frac{2}{\pi^{2}}x^{2},\qquad x\in(0,\pi),$
which means that the previous integral expression is bounded from above by
$\frac{\pi}{\sqrt{2}}\frac{1}{\sqrt{\sinh(t-t^{\prime})\,\sinh(t^{\prime})}}\,\int_{\theta_{0}}^{\frac{\pi}{2}}\frac{d\theta}{\theta}\leq\frac{\pi^{2}}{2\sqrt{2}}\,\frac{1}{\sqrt{\sinh(t-t^{\prime})\,\sinh(t^{\prime})}}\,\frac{1}{\theta_{0}}.$
Let us bound the third term of Expression (7.5). Since $t-t^{\prime}>\rho$ and $t^{\prime}>r$, we have the following upper bound for it:
$\int_{\frac{\pi}{2}}^{\pi}\frac{d\theta}{\sqrt{\sinh(t-t^{\prime})\,\sinh(t^{\prime})}}=\frac{\pi}{2}\,\frac{1}{\sqrt{\sinh(t-t^{\prime})\,\sinh(t^{\prime})}}$
We now have an upper bound for Expression (7.5) and thus an upper bound for Expression (7.2), the latter being:
$2\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{\cosh(t^{\prime})-\cosh(r)}\frac{\theta_{0}}{\sqrt{\cosh(t-t^{\prime})}}+\frac{\pi^{2}}{\sqrt{2}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{\sqrt{\cosh(t^{\prime})-\cosh(r)}}\,\frac{1}{\sqrt{\sinh(t-t^{\prime})\,\sinh(t^{\prime})}}\,\frac{1}{\theta_{0}}$
$+\ \pi\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{\sqrt{\cosh(t^{\prime})-\cosh(r)}\sqrt{\sinh(t-t^{\prime})\sinh(t^{\prime})}}$
We choose
@@
$\theta_{0}=\min\left(\frac{\pi}{2},\left(\frac{\cosh(t^{\prime})-\cosh(r)}{\sinh(t^{\prime})}\right)^{\alpha}\right),$
with $\alpha\in(0,\frac{1}{2})$. The previous integral expression has the following upper bound:
$2\,\frac{1}{\sqrt{\cosh(t-t^{\prime})}\,\sinh(t^{\prime})^{\alpha}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{(\cosh(t^{\prime})-\cosh(r))^{1-\alpha}}$
$+\ \frac{\pi^{2}}{\sqrt{2}}\,\frac{1}{\sqrt{\sinh(t-t^{\prime})}\sinh(t^{\prime})^{\frac{1}{2}-\alpha}}\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{(\cosh(t^{\prime})-\cosh(r))^{\frac{1}{2}+\alpha}}$
$+\ \pi\int_{0}^{t^{\prime}}\frac{\sinh(r)\,dr}{\sqrt{\cosh(t^{\prime})-\cosh(r)}\sqrt{\sinh(t-t^{\prime})\sinh(t^{\prime})}}$
The upper bound was obtained by replacing $\theta_{0}$ with
$\left(\frac{\cosh(t^{\prime})-\cosh(r)}{\sinh(t^{\prime})}\right)^{\alpha},$
which is possible since the middle term would vanish if the previous ratio is larger than $\frac{\pi}{2}$. The previous integral expression has the following equivalent expression obtained by integrating:
$2\,\frac{1}{\sqrt{\cosh(t-t^{\prime})}\,\sinh(t^{\prime})^{\alpha}}\,\frac{(\cosh(t^{\prime})-1)^{\alpha}}{\alpha}+\frac{\pi^{2}}{\sqrt{2}}\,\frac{1}{\sqrt{\sinh(t-t^{\prime})}\sinh(t^{\prime})^{\frac{1}{2}-\alpha}}\frac{(\cosh(t^{\prime})-1)^{\frac{1}{2}-\alpha}}{\frac{1}{2}-\alpha}$
$+\ \pi\frac{\sqrt{\cosh(t^{\prime})-1}}{\sqrt{\sinh(t^{\prime})}}\,\frac{1}{\sqrt{\sinh(t-t^{\prime})}}.$
By choosing $\alpha=\frac{1}{4}$, we have the following upper bound for the previous expression:
$40\,\frac{1}{\sqrt{\sinh(t-t^{\prime})}}.$
This is also the desired upper bound for $F_{t,t^{\prime},\rho}$.
We state the second lemma concerning the case of $\rho\in(t-t^{\prime},t)$.
@@
Let
where $z,z^{\prime}\in\mathbb{H}$ are any two points separated by distance $\rho$. Assume that $t>t^{\prime}>0$. Assume also that $\rho\in(t-t^{\prime},t)$. Then
$F_{t,t^{\prime},\rho}\leq\frac{C_{3}}{\sqrt{\sinh(t-t^{\prime})}},$
where $C_{3}=50$.
The following figure, Figure 3, illustrates the integration domain.
Figure 3. The case of $\rho\in(t-t^{\prime},t)$. The integration domain $B(z,t)\cap B(z^{\prime},t^{\prime})$ is the intersection of two balls. The center $z^{\prime}$ of the smaller ball is contained in the larger ball. The integration domain is split into two parts, $D_{1}$ and $D_{2}$, by two geodesics of length $t^{\prime}$ going from $z^{\prime}$ to the points $w$ and $w^{\prime}$ where the two balls intersect.
In this case, the integration domain is not a ball, since $\partial B(z,t)$ and $\partial B(z^{\prime},t^{\prime})$ intersect; call the intersections $w$ and $w^{\prime}$. The segments $z^{\prime},w]$ and $z^{\prime},w^{\prime}]$ split the integration domain into two disjoint parts, which are treated separately in the proof. For the other side ($D_{1}$ in Figure 3), the idea is to use a similar approach as in the proof of Lemma 7.4: use polar coordinates around $z^{\prime}$ and deal with the small and large angles separately. The remaining side ($D_{2}$ in Figure 3) is easier and can be handled using polar coordinates around $z$.
@@
By our assumption, we know $\partial B(z,t)$ and $\partial B(z^{\prime},t^{\prime})$ intersect at exactly two points. Call these points $w$ and $w^{\prime}$. The segments $z^{\prime},w]$ and $z^{\prime},w^{\prime}]$ split the integration domain $B(z,t)\cap B(z^{\prime},t^{\prime})$ into two disjoint parts, $D_{1}$ and $D_{2}$.
Hence $F_{t,t^{\prime},\rho}$ can be written as
$\int_{D_{1}}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(x,z))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(x,z^{\prime}))}}$
$+\ \int_{D_{2}}\frac{dx}{\sqrt{\cosh(t)-\cosh(d(x,z))}\,\sqrt{\cosh(t^{\prime})-\cosh(d(x,z^{\prime}))}}.$ (7.6)
We bound the two terms separately.
We begin by bounding the integration over $D_{1}$. We write $x$ using polar coordinates around $z^{\prime}$, taking $\theta=0$ as the angle in the direction $z,z^{\prime}](z^{\prime})$. Hence $x=\pi_{1}(\varphi_{r}(z^{\prime},\theta))$. $R(r,\theta)$ is $d(x,z)$ in the polar coordinates. We notice that the integration domain of the angles is $(\Theta,2\pi-\Theta)$, where $\Theta$ is the angle between $z,z^{\prime}](z^{\prime})$ and $z^{\prime},w](z^{\prime})$. The following figure, Figure 4, illustrates this.
Figure 4. The integration domain $D_{1}$ consists of the points that are outside the sector formed by $w,z^{\prime},$ and $w^{\prime}$ in ball $B(z^{\prime},t^{\prime})$. Polar coordinates around $z^{\prime}$ are used.
Hence the first term of Expression (7.6) equals